OpenStudy (anonymous):

A hare, pursued by a grey-hound is 24 of her own leaps ahead of him. While the hare takes 6 leaps, the grey-hound takes 7 leaps. 7 leaps of grey-hound is equal to 9 leaps of hare. In how many leaps will the grey-hound overtake the hare?

imqwerty (imqwerty):

ok :) let us assume that one leap of hare=1m so we are given that 7 leaps of hound=9 leaps of hare so 1leap of hound=9/7 leaps of hare so 1leap of hound=9/7m nd then lets assume that the hare makes one leap in 1 sec time taken by hare to take 6 leaps =6secs and 6sec=time taken by hound to make 7 leaps so the hound makes 7/6leaps per sec 1 leap of hound=9/7m so in 1 sec hound travels 7/6 x 9/7 m=3/2m speed of hare is 1m/sec and speed of hound is 1.5m/sec and hare is 24 of her own leaps ahead =24m so let the time taken by hound to capture the hare be t so distance travelled by hound=distance traveled by hare+24m 1.5(t) = 1(t)+24 0.5(t)=24 t=24/0.5 t=48secs hound makes 7/6 leaps per sec so leaps made in 48 secs=7/6 x 48 =56

OpenStudy (anonymous):

The given answer is 280 :/

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

:) lets recheck

OpenStudy (anonymous):

6sec=time taken by hound to make 7 leaps Can you explain this?

imqwerty (imqwerty):

its given that when the hare makes leaps the hound makes 7 leaps so the time taken by hare to make 6 leaps=6sec=time taken by hound to make 7 leaps

OpenStudy (anonymous):

Okay. But we ain't getting the right answer.

OpenStudy (anonymous):

@imqwerty ?

imqwerty (imqwerty):

umm...m searching for a mistake in my method..